### v1

# High-order WENO finite-difference methods for hyperbolic nonconservative systems of Partial Differential Equations

Abstract
This work aims to extend the well-known high-order WENO finite-difference methods for systems of conservation laws to nonconservative hyperbolic systems. The main difficulty of these systems both from the theoretical and the numerical points of view comes from the fact that the definition of weak solution is not unique: according to the theory developed by Dal Maso, LeFloch, and Murat in 1995, it depends on the choice of a family of paths. A general strategy is proposed here in which WENO operators are not only used to reconstruct fluxes but also the nonconservative products of the system. Moreover, if a Roe linearization is available, the nonconservative products can be computed through matrix-vector operations instead of path-integrals. The methods are extended to problems with source terms and two different strategies are introduced to obtain well-balanced schemes. These numerical schemes will be then applied to the two-layer shallow water equations in one- and two- dimensions to obtain high-order methods that preserve water-at-rest steady states.
Keywords: WENO finite difference scheme, High order accuracy, Well-balanced scheme, Nonconservative equations, Path-conservative method.
1 Introduction
We aim to construct well-balanced high-order WENO finite-difference schemes for hyperbolic nonconservative problems of the form
$U_{t}+A_{1}(U)U_{x}+A_{2}(U)U_{y}=S_{1}(U)H_{x}+S_{2}(U)H_{y},$ (1)
where the unknown $U(x,y,t)$ takes value in an open convex set $\Omega\in\mathbb{R}^{N}$; $A_{i}(U)$, $i=1,2$ are smooth matrix-valued functions; $S_{i}(U)$, $i=1,2$ are vector-valued functions, and $H(x,y)$ is a known function from $\mathbb{R}^{2}$ to $\mathbb{R}$. The 1D case
$U_{t}+A(U)U_{x}=S(U)H_{x}$ (2)
will be also considered.
PDE systems of the form
$U_{t}+F_{1}(U)_{x}+F_{2}(U)_{y}+B_{1}(U)U_{x}+B_{2}(U)U_{y}=S_{1}(U)H_{x}+S_{2}(U)H_{y},$ (3)
where $F_{i}(U)$, $i=1,2$ are the flux function and $B_{i}(U)$, $i=1,2$ are matrix-valued functions, can be considered as particular cases of (1). Systems of conservation laws ($B_{i}\equiv 0$, $S_{i}\equiv 0$, $i=1,2)$ and systems of balance laws ($B_{i}\equiv 0$, $i=1,2)$ are in turn particular cases of (3).
The major difficulty of systems (1) or (3), from both the theoretical and the numerical points of view, comes from the fact that the presence of nonconservative products makes that, unlike for systems of conservation laws, the definition of weak solution is not unique. In the theory developed by Dal Maso, LeFloch, and Murat in [35], nonconservative products are defined as Borel measures based on the choice of a family of paths, i.e. a Lipschitz-continuous function $\Psi:[0,1]\times\Omega\times\Omega\to\Omega$ satisfying
$\Psi(0;U_{L},U_{R},0)=U_{L},\quad\Psi(1;U_{L},U_{R})=U_{R}$
for all $U_{L}$, $U_{R}\in\Omega$, and
$\Psi(s;U,U)=U$
for all $U\in\Omega$ and $s\in[0,1]$. Once the family of paths has been used, the generalized Rankine-Hugoniot conditions satisfied by the admissible weak solutions at a jump are the following
$\sigma(U^{+}-U^{-})=\int_{0}^{1}\left(\sum_{i=1}^{2}n_{i}A_{i}(\Psi(s;U^{-},U^{+}))\right)\partial_{s}\Psi(s;U^{-},U^{+})\,ds,$
where $\sigma$ is the propagation speed; $U^{\pm}$ the lateral limits of the solution; and $\vec{n}=(n_{1},n_{2})$ a unit vector normal to the jump.
Based on this theory, a framework to design finite-volume methods for nonconservative sysmes was introduced in [37] based on the concept of path-conservative methods. These methods have been extensively applied to solve nonconservative systems: see, for instance, [8, 11, 13, 16, 17, 19].
In this paper, we focus on WENO finite-difference schemes. In the past decades, these methods have been widely applied to systems of conservation and balance laws: see for instance [5, 7, 28, 29, 43, 44, 48]. In [22] a fifth-order A-WENO finite-difference scheme for 1D and 2D systems of nonconservative hyperbolic systems was introduced. This scheme is based on the path-conservative central-upwind method and the global flux approach: the integral of the nonconservative term is considered as a new flux function of the system.
The goal of this paper is to introduce a new approach to design high-order finite-difference methods for nonconservative systems. The key idea is to apply a standard high-order WENO reconstruction operator to the nonconservative products computed using the selected family of paths: more precisely, in the 1D case, the selected WENO operator will be applied to reconstruct quantities of the form
$\int_{0}^{1}A^{\pm}(\Psi(s;U_{i},U_{j}))\partial_{s}\Psi(s;U_{i},U_{j})\,ds,\quad j\in\mathcal{S}_{i},$
where $\mathcal{S}_{i}$ represents the stencil of the $i$th point and $A^{\pm}(U)$ represents a splitting of the matrix system $A(U)$. Two different splittings will be considered here: Lax-Friedrichs and upwind splittings. It will be seen that, if a path-consistent Roe linearization is available (see [15]), the quantities to be reconstructed can be computed by using matrix-vector products instead of path-integrals. Compared to the approach in [22], the main advantages of this new one are the following:
  • the accuracy in space of the methods only depends on the order of the standard WENO reconstruction selected: no order-dependent corrections of the fluxes are necessary;
  • no integrals involving quadrature points in the cells have to be computed which avoids having to use an additional reconstruction operator with uniform accuracy in the entire cells.
The numerical methods will be applied to the 1D and 2D hyperbolic two-layer shallow-water equations that govern the flow of two superposed layers of immiscible homogeneous fluids. This system is used in different ocean and coastal engineering simulations and there is a vast literature focusing on its numerical analysis: see for instance [10, 12, 21, 22, 23, 33, 36].
A relevant property to be satisfied by the numerical methods solving systems of the form (1) or (3) is the preservation of some or all the steady-state solutions of the system, i.e. the well-balanced property. For instance, in the context of the one or the two-layer shallow-water equations, a minimal requirement to the numerical methods is to exactly preserve the steady states corresponding to water-at-rest, i.e. to satisfy the C-property according to [4]. Different techniques have been proposed to design well-balanced schemes including hydrostatic reconstruction related [2, 18], relaxation methods [1], consistent discretization of the flux and source term [32, 46, 47], etc. See also [6, 26, 34].
In the context of finite-volume methods, a general strategy to design well-balanced methods has been described in [20]. In this strategy, a stationary solution whose average is the numerical approximation at every cell has to be computed at every time step. Then, a standard reconstruction operator is applied to the differences of the cell values at the stencil and the cell averages of the local stationary solution. In [38] this strategy has been extended to WENO finite-difference methods for systems of balance laws. Two different strategies will be followed here to obtain well-balanced numerical methods: one of them is the extension to nonconservative systems of the strategy introduced in this last reference, while the other consists in combining the upwind splitting with an adequate choice of family of paths.
As it is well known (see [14]), in the case of nonconservative systems, the numerical solutions obtained with finite-difference or similar methods that are formally consistent with the definition of weak solution related to a given family of paths may converge to functions that are not weak solutions according to that family. Nevertheless, it will be seen in Section 6 that the numerical results obtained for the two-layer shallow-water equations are similar to those obtained with other methods. Nevertheless, in order to ensure the convergence to functions that are weak solutions according to the selected family of paths, the numerical dissipation close to shocks has to be controlled: see for instance [3]. In [39, 40] high-order finite-volume numerical methods that are able to correctly capture isolated shock waves have been designed based on the use of a discontinuous reconstruction operator in cells where a shock is detected: similar strategies can be adapted to the numerical methods introduced here, what will be done in future works.
The rest of the paper is organized as follows. In Section 2, firstly, the new WENO path-conservative schemes for 1D homogeneous (i.e. without source terms) nonconservative systems are introduced and their high-order accuracy property is proved. In Section 3, source terms are included and two strategies to obtain well-balanced methods are described. In Section 4, the proposed schemes are extended to 2D nonconservative systems. In Section 5, we apply the proposed scheme to 1D and 2D two-layer shallow water equations: the numerical results are presented in Section 6. Finally, some conclusions are drawn in Section 7.
2 Path-Conservative WENO finite-difference reconstruction methods
2.1 WENO methods
The goal of this paper is to extend high-order finite-difference methods based on flux reconstructions for conservation laws system
$U_{t}+F(U)_{x}=0,$ (4)
to nonconservative systems
$U_{t}+A(U)U_{x}=0.$ (5)
It will be assumed here that the system is strictily hyperbolic, i.e. for every $U$, the matrix $A(U)$ has N different real eigenvalues
$\lambda_{1}(U),\dots,\lambda_{N}(U).$
In particular, systems of the form
$U_{t}+F(U)_{x}+B(U)U_{x}=0$ (6)
will be considered that can be written in the form (5) with
$A(U)=J(F(U))+B(U),$
where $J(F(U))$ represents the Jacobian of the flux function $F(U)$.
Semi-discrete high-order finite-difference methods for systems of conservation laws (4) have the form:
$\frac{dU_{i}}{dt}+\frac{1}{\Delta x}\left({F}_{i+1/2}-{F}_{i-1/2}\right)=0,$ (7)
where $F_{i+1/2}$ is a high-order reconstruction of the flux function. Uniform meshes of constant step $\Delta x$ will be considered and the following notation will be also used for the intercells
$x_{i+1/2}=x_{i}+\frac{\Delta x}{2}.$
In the particular case of the WENO reconstruction of order $p=2k+1$ for systems of balance laws, two flux reconstructions are computed using the values at the points $x_{i-k},\dots,x_{i+k}$:
$F^{L}_{i+1/2}$ $=$ $\mathcal{R}^{L}(F(U_{i-k}),\dots,F(U_{i+k})),$ (8)
$F^{R}_{i-1/2}$ $=$ $\mathcal{R}^{R}(F(U_{i-k}),\dots,F(U_{i+k})).$ (9)
These are the so-called left- and right-biased reconstructions, related throught the equality:
$\mathcal{R}^{L}(F(U_{i-k}),\dots,F(U_{i+k}))=\mathcal{R}^{R}(F(U_{i+k}),\dots,F(U_{i-k})).$
In order to compute the numerical flux $F_{i+1/2}$, first a splitting of the flux function is considered
$F(U)=F^{+}(U)+F^{-}(U),$
in such a way that the eigenvalues of the Jacobian $J^{+}(U)$ (resp. $J^{-}(U)$) of $F^{+}(U)$ (resp. $F^{-}(U)$) are positive (resp. negative). A standard choice is the Lax-Friedrichs flux-splitting:
$F^{\pm}(U)=\frac{1}{2}\left(F(U)\pm\alpha U\right),$
where $\alpha$ is the maximum of the absolute value of the eigenvalues of $\{J(U_{i})\}$, this maximum being taken over either local (WENO-LLF) or global (WENO-LF): see [29, 43].
Then, the reconstruction operator is applied to $F^{\pm}$:
$F^{+}_{i+1/2}$ $=$ $\mathcal{R}^{L}(F^{+}(U_{i-k}),\dots,F^{+}(U_{i+k})),$ (10)
$F^{-}_{i+1/2}$ $=$ $\mathcal{R}^{R}(F^{-}(U_{i+1-k}),\dots,F^{-}(U_{i+1+k})),$ (11)
and finally,
$F_{i+1/2}=F^{+}_{i+1/2}+F^{-}_{i+1/2}.$ (12)
The reconstruction then satisfies
$\frac{1}{\Delta x}\left(F_{i+1/2}-F_{i-1/2}\right)=F(U)_{x}+O(\Delta x^{2k+1}),\quad\forall i.$
In order to extend these numerical methods to (5), let us first rewrite (7) as follows:
$\frac{dU_{i}}{dt}+\frac{1}{\Delta x}\left({F}_{i+1/2}-{F}\left(U_{i}\right)+{F}\left(U_{i}\right)-{F}_{i-1/2}\right)=0,$ (13)
or, equivalently
$U_{i}^{\prime}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=0,$ (14)
with
$\widehat{D}^{-}_{i+1/2}={F}_{i+1/2}-{F}\left(U_{i}\right),\quad\widehat{D}^{+}_{i-1/2}={F}\left(U_{i}\right)-{F}_{i-1/2}.$
One has then
$\widehat{D}^{-}_{i+1/2}$ $=$ ${F}_{i+1/2}-{F}(U_{i})$
$=$ $F^{+}_{i+1/2}-{F^{+}}(U_{i})+F^{-}_{i+1/2}-{F^{-}}(U_{i})$
$=$ $\mathcal{R}^{L}(F^{+}(U_{i-k}),\dots,F^{+}(U_{i+k}))-{F^{+}}(U_{i})$
$+$ $\mathcal{R}^{R}(F^{-}(U_{i+1-k}),\dots,F^{-}(U_{i+1+k}))-{F^{-}}(U_{i})$
$=$ $\mathcal{R}^{L}(F^{+}(U_{i-k})-{F^{+}}(U_{i}),\dots,F^{+}(U_{i+k})-{F^{+}}(U_{i}))$
$+$ $\mathcal{R}^{R}(F^{-}(U_{i+1-k})-{F^{-}}(U_{i}),\dots,F^{-}(U_{i+1+k})-{F^{-}}(U_{i}))$
$=$ $\mathcal{R}^{L}(D^{+}_{i,i-k},\dots,D^{+}_{i,i+k})+\mathcal{R}^{R}(D^{-}_{i,i+1-k},\dots,D^{-}_{i,i+1+k}),$
where the following notation has been used
$D^{\pm}_{j,k}=F^{\pm}(U_{k})-F^{\pm}(U_{j}),\quad\forall j,k.$
Analogously
$\widehat{D}^{+}_{i-1/2}$ $=$ ${F}(U_{i})-F_{i-1/2}$
$=$ $\mathcal{R}^{L}(D^{+}_{i-1-k,i},\dots,D^{+}_{i-1+k,i})+\mathcal{R}^{R}(D^{-}_{i-k,i},\dots,D^{-}_{i+k,i}).$
In order to extend the method to (5) first a splitting of $A(U)$
$A(U)=A^{+}(U)+A^{-}(U),$
has to be chosen in such a way that the eigenvalues of $A^{+}(U)$ (resp. $A^{-}(U)$) are positive (resp. negative). For instance, the LF splitting can be extended in a natural manner:
$A^{\pm}(U)=\frac{1}{2}\left(A(U)\pm\alpha I\right),$
with $\alpha$ is computed as in the conservative case.
Finally, the last ingredient required to extend the numerical methods to nonconservative systems is a family of paths $\Psi:[0,1]\times\Omega\times\Omega\to\Omega$. Using $\Psi$ we have:
$D^{\pm}_{j,k}=F^{\pm}(U_{k})-F^{\pm}(U_{j})=\int_{0}^{1}DF^{\pm}(\Psi(s;U_{j},U_{k}))\partial_{s}\Psi(s;U_{j},U_{k})\,ds.$
The natural extension of the numerical method (7) to the system (5) is then given by (14) with
$\widehat{D}^{-}_{i+1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i,i-k},\dots,D^{+}_{i,i+k})+\mathcal{R}^{R}(D^{-}_{i,i+1-k},\dots,D^{-}_{i,i+1+k}),$ (15)
$\widehat{D}^{+}_{i-1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i-1-k,i},\dots,D^{+}_{i-1+k,i})+\mathcal{R}^{R}(D^{-}_{i-k,i},\dots,D^{-}_{i+k,i}).$ (16)
and
$D^{\pm}_{j,k}=\int_{0}^{1}A^{\pm}(\Psi(s;U_{j},U_{k}))\partial_{s}\Psi(s;U_{j},U_{k})\,ds.$ (17)
Observe that, while in the case of a system of conservation laws, the resulting numerical method is independent of the chosen family of paths (since it is equivalent to (7)), for nonconservative systems the numerical method depends on the chosen family of paths. Please note that, while for systems of conservation laws only two reconstructions per intercell are required, here 4 reconstructions are needed in $x_{i+1/2}$: two reconsttructions to compute $\widehat{D}^{+}_{i+1/2}$ and two others to compute $\widehat{D}^{-}_{i+1/2}$
Let us suppose that a Roe linearization is available for the chosen family of paths $\Psi$, i.e. a matrix-valued function $A_{\Psi}:\Omega\times\Omega\mapsto\mathbb{R}^{N}\times\mathbb{R}^{N}$ that satisfies the following properties:
  1. 1.
    For each $U,V\in\Omega$ , $A_{\Psi}\left(U,V\right)$ has $N$ distinct real eigenvalues:
    $\lambda_{1}\left(U,V\right)<\lambda_{2}\left(U,V\right)<\cdots<\lambda_{N}\left(U,V\right).$
  2. 2.
    $A_{\Psi}(U,U)=A(U)$, for every $U\in\Omega$.
  3. 3.
    For any $U,V\in\Omega$,
[…]
Then, the fluctuations $D_{j,k}$ can be computed as follows:
$D^{\pm}_{j,k}=A^{\pm}_{\Psi}\left(U_{j},U_{k}\right)\left(U_{k}-U_{j}\right),$ (19)
where $A^{\pm}_{\Psi}\left(U,V\right)$ represents the application to $A_{\Psi}\left(U,V\right)$ of the splitting chosen for $A(U)$: for instance, if the LF splitting has been chosen, then
$A^{\pm}_{\Psi}\left(U,V\right)=\frac{1}{2}\left(A_{\Psi}\left(U,V\right)\pm\alpha I\right).$
It is well-known fact that high-order finite-difference methods are less oscillatory if the reconstructions are performed in characteristic fields. To do that, let us suppose that a Roe linearization $A_{\Psi}$ is available. Given two indices $j,k$ the following notation will be used
$A_{j,k}=A_{\Psi}(U_{j},U_{k})$
to represent the Roe matrix. This matrix can be decomposed in the form:
$A_{j,k}={L}_{j,k}\Lambda_{j,k}{R}_{j,k},$
where $\Lambda_{j,k}$ is the $N$-dimensional diagonal matrix whose coefficients are the eigenvalues of $A_{j,k}$:
$\lambda_{j,k;1},\dots,\lambda_{j,k;N};$
$R_{j,k}$ is a matrix whose $l$th column $\vec{r}_{j,k;l}$ is an eigenvector associated to $\lambda_{j,k;l}$; and ${L}_{j,k}={R_{j,k}^{-1}}$ is a matrix whose arrows are left-eigenvalues. To compute the reconstruction in characteristic variables, the fluctuations involved in the computation of $\widehat{D}^{\pm}_{i+1/2}$ are first multiplied by the matrix $L_{i,i+1}$ of left eigenvectors of $A_{i,i+1}$ and the reconstructed fluxes are then multiplied by the matrix $R_{i,i+1}$ of right eigenvectors, i.e.
$\widehat{D}^{-}_{i+1/2}$ $=$ $R_{i,i+1}\mathcal{R}^{L}(L_{i,i+1}D^{+}_{i,i-k},\dots,L_{i,i+1}D^{+}_{i,i+k})$ (20)
$+R_{i,i+1}\mathcal{R}^{R}(L_{i,i+1}D^{-}_{i,i+1-k},\dots,L_{i,i+1}D^{-}_{i,i+1+k}),$
$\widehat{D}^{+}_{i-1/2}$ $=$ $R_{i-1,i}\mathcal{R}^{L}(L_{i-1,i}D^{+}_{i-1-k,i},\dots,L_{i-1,i}D^{+}_{i-1+k,i})$ (21)
$+R_{i-1,i}\mathcal{R}^{R}(L_{i-1,i}D^{-}_{i-k,i},\dots,L_{i-1,i}D^{-}_{i+k,i}).$
An alternative implementation of WENO methods is given by the upwind approach, in which the splitting of the Roe matrix is given by:
$A^{\pm}_{\Psi}\left(U,V\right)=P^{\pm}_{\Psi}\left(U,V\right)A_{\Psi}\left(U,V\right).$
Here, $P^{\pm}$ represent the projection matrices onto the positive/negative eigenvectors of the Roe matrix, i.e.
$P^{\pm}_{\Psi}\left(U,V\right)=R_{\Psi}\left(U,V\right)M^{\pm}_{\Psi}\left(U,V\right)R^{-1}_{\Psi}\left(U,V\right),$
where $M^{\pm}_{\Psi}\left(U,V\right)$ is the diagonal matrix whose coefficients are
$\frac{1}{2}\left(1\pm\operatorname{sign}\left(\lambda_{j}(U,V)\right)\right),\quad j=1,\ldots,N,$
and $R_{\Psi}\left(U,V\right)$ is a matrix whose $j$th columns is an eigenvector associated to $\lambda_{j}(U,V)$. This strategy is implemented as follows: given two indices $j,k$, let $\{\alpha_{j,k;l}\}_{l=1}^{N}$ be the coordinates of $U_{k}-U_{j}$ in the basis of eigenvectors of the Roe matrix $A_{j,k}$, i.e.
$U_{k}-U_{j}=\sum_{l=1}^{N}\alpha_{j,k;l}\vec{r}_{j,k;l}.$
Then
$D^{\pm}_{j,k}=A^{\pm}_{j,k}(U_{k}-U_{j})=\sum_{l=1}^{N}\alpha_{j,k;l}\lambda^{\pm}_{j,k;l}\vec{r}_{j,k;l},$
where, given $\lambda\in\mathbb{R}$, $\lambda^{\pm}$ represent the positive and negative part of $\lambda$, i.e.
$\lambda^{+}=\frac{\lambda+|\lambda|}{2},\quad\lambda^{-}=\frac{\lambda-|\lambda|}{2}.$
Observe that, if the LF splitting is chosen and reconstruction in characteristic variables is performed, the right and left eigenvectors of every Roe matrix have to be computed. On the other hand, if the upwind approach is followed, the eigenvectors and eigenvalues of the Roe matrices $A_{k,j}$ are required. Moreover, the linear system
$R_{j,k}\vec{\alpha}_{j,k}=U_{k}-U_{j},$
has to be solved to compute the coordinates $\alpha_{j,k;l}$. Therefore, it is more computationally expensive.
Remark.
Since the expression of WENO reconstructions is a linear combination of the fluxes whose coefficients depend nonlinearly on the data through the smoothness indicators, it can be shown that, for problems of the form (6), the numerical method (7) can be written in the form
$U_{i}^{\prime}=-\frac{1}{\Delta x}\left(F_{i+1/2}-F_{i-1/2}+\widehat{B}^{-}_{i+1/2}+\widehat{B}^{+}_{i-1/2}\right),$ (22)
where $F_{i+1/2}$ and $\widehat{B}_{i+1/2}$ are, respecitvely, standard WENO reconstructions of the flux function and the nonconservative terms, in which the nonlinear coefficients are the same.
2.2 Accuracy of the methods
Let us check that (14)-(15)-(16)-(17) is a high-order numerical method for (5).
Proposition 1.
Let us consider a smooth solution $U(x,t)$ of (5) and assume that $A(U)$ and $\Psi$ are smooth. We also assume that $\Psi$ satisfies
$\int_{0}^{1}A(\Psi(s;V,U)\partial_{s}\Psi(s;V,U)\,ds=-\int_{0}^{1}A(\Psi(s;U,V)\partial_{s}\Psi(s;U,V)\,ds$
for all $U,V\in\Omega$. Then we have
$\partial_{t}U(x_{i},t)+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=O(\Delta x^{2k+1}),$ (23)
where $p=2k+1$ is the order of the reconstruction operator.
Proof.
Given an index $i$ and a time $t$, let us define the function $G^{t}_{i}(x)$ as follows:
$G^{t}_{i}(x)=\int_{0}^{1}A(\Psi(s;U(x_{i},t),U(x,t))\partial_{s}\Psi(s;U(x_{i},t),U(x,t))\,ds.$
This function satisfies
$G^{t}_{i}(x_{i})=0,\quad\partial_{x}G_{i}^{t}(x_{i})=A(U(x_{i},t))U_{x}(x_{i},t).$
In effect,
$G_{i}^{t}(x_{i})=\int_{0}^{1}A(\Psi(s;U(x_{i},t),U(x_{i},t))\partial_{s}\Psi(s;U(x_{i},t),U(x_{i},t))\,ds=0,$
since
$\Psi(s;U(x_{i},t),U(x_{i},t))=U(x_{i},t),\quad\forall s.$
On the other hand:
$\partial_{x}G^{t}_{i}(x_{i})$ $=$ $\lim_{h\to 0}\frac{G_{i}^{t}(x_{i}+h)-G^{t}_{i}(x_{i})}{h}$
$=$ $\lim_{h\to 0}\frac{1}{h}\int_{0}^{1}A(\Psi(s;U(x_{i},t),U(x_{i}+h,t)))\partial_{s}\Psi(s;U(x_{i},t),U(x_{i}+h,t))\,ds$
$=$ $\lim_{h\to 0}\int_{0}^{1}\frac{1}{h}\left(A(\Psi(s;U(x_{i},t),U(x_{i}+h,t)))-A(U(x_{i},t))\right)\partial_{s}\Psi(s;U(x_{i},t),U(x_{i}+h,t))\,ds$
$\quad+\lim_{h\to 0}A(U(x_{i},t))\frac{1}{h}\int_{0}^{1}\partial_{s}\Psi(s;U(x_{i},t),U(x_{i}+h,t))\,ds$
$=$ $\lim_{h\to 0}\int_{0}^{1}\frac{1}{h}\left(A(\Psi(s;U(x_{i},t),U(x_{i}+h,t)))-A(U(x_{i},t))\right)\partial_{s}\Psi(s;U(x_{i},t),U(x_{i}+h,t))\,ds$
$\quad+\lim_{h\to 0}A(U(x_{i},t))\frac{U(x_{i}+h,t)-U(x_{i},t)}{h}$
$=$ $A(U(x_{i},t))U_{x}(x_{i},t),$
where, in the first term, it has been used again that
$\partial_{s}\Psi(s;U(x_{i},t),U(x_{i}+h,t))\to\partial_{s}\Psi(s;U(x_{i},t),U(x_{i},t))=0.$
Observe that, for all $j$:
$D_{i,j}=G^{t}_{i}(x_{j}),\quad D_{j,i}=-G_{i}^{t}(U)(x_{j}).$
Therefore
$\widehat{D}^{-}_{i+1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i,i-k},\dots,D^{+}_{i,i+k})+\mathcal{R}^{R}(D^{-}_{i,i+1-k},\dots,D^{-}_{i,i+1+k})$
$=$ $\mathcal{R}^{L}(G_{i}^{t,+}(x_{i-k}),\dots,G_{i}^{t,+}(x_{i+k}))+\mathcal{R}^{R}(G^{t,-}_{i}(x_{i+1-k}),\dots,G^{t,-}_{i}(x_{i+1+k}))=\hat{G}^{t}_{i,i+1/2},$
$\widehat{D}^{+}_{i-1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i-1-k,i},\dots,D^{+}_{i-1+k,i})+\mathcal{R}^{R}(D^{-}_{i-k,i},\dots,D^{-}_{i+k,i})$
$=$ $-\mathcal{R}^{L}(G^{t,+}_{i}(x_{i-1-k}),\dots,G^{t,+}_{i}(x_{i-1+k}))-\mathcal{R}^{R}(G^{t,-}_{i}(x_{i-k}),\dots,G^{t,-}_{i}(x_{i+k}))=-\hat{G}_{i,i-1/2},$
where $G_{i}^{t,\pm}$ represents the splitting of the function $G_{i}^{t}$ and $\hat{G}^{t}_{i,i\pm 1/2}$ is its WENO reconstruction. Therefore we have:
$\frac{1}{\Delta x}(\widehat{D}_{i-1/2}^{+}+\widehat{D}_{i+1/2}^{-})$ $=$ $\frac{1}{\Delta x}(\hat{G}_{i,i+1/2}-\hat{G}_{i,i-1/2})$
$=$ $\partial_{x}G^{t}_{i}(x_{i})+O(\Delta x^{2k+1})$
$=$ $A(U(x_{i},t))U_{x}(x_{i},t)+O(\Delta x^{2k+1}),$
which leads to (23). ∎
3 Problems with source terms and well-balanced property
Let us consider now problems with source term
$U_{t}+A(U)U_{x}=S(U)H_{x},$ (24)
where $H(x)$ is a known function. The well-balanced property of the methods is related to the preservation of the stationary solutions $U^{*}$ of the system, which satisfy the equations
$A(U^{*})U^{*}_{x}=S(U^{*})H_{x}.$
Two strategies are considered here to discretize the source term:
3.1 Strategy 1
The first strategy consists in writing (24) in the form (5) as follows:
$W_{t}+\widetilde{A}(W)W_{x}=0,$
with
$W=\left[\begin{array}[]{c}U\\ H\end{array}\right]\in\Omega\times\mathbb{R},\quad\mathcal{A}(W)=\left[\begin{array}[]{c|c}A(U)&-S(U)\\ \hline\cr 0&0\end{array}\right],$
and then the general strategy is applied to derive high-order WENO finite-difference methods. To do this, a family of paths
$\widetilde{\Psi}(s;W_{L},W_{R})=\left[\begin{array}[]{c}\Psi_{U}(s;W_{L},W_{R})\\ \Psi_{H}(s;W_{L},W_{R})\end{array}\right]$
has to be chosen. If the artificial unknown $H$ is dropped, the numerical method writes again as (14)-(15)-(16) where now
$D_{j,k}=\int_{0}^{1}A(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{U}(s;W_{j},W_{k})\,ds-\int_{0}^{1}S(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{H}(s;W_{j},W_{k})\,ds,$ (25)
where the notation $W_{j}=[U_{j},H(x_{j})]^{T}$ has been used. Moreover, if a Roe linearization is available, i.e. if a matrix-valued function $A_{\widetilde{\Psi}}(W_{L},W_{R})$ and a vector-valued function $S_{\widetilde{\Psi}}(W_{L},W_{R})$ exist such that
  • $A_{\widetilde{\Psi}}(W_{L},W_{R})$ has real different eigenvalues;
  • $A_{\widetilde{\Psi}}(W,W)=A(U)$ for all $W=[U,H]^{T}$;
  • for all $W_{L},W_{R}\in\Omega\times\mathbb{R}$
    $A_{\widetilde{\Psi}}(W_{L},W_{R})(U_{R}-U_{L})-S_{\widetilde{\Psi}}(W_{L},W_{R})(H_{R}-H_{L})$
    $=\int_{0}^{1}A(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{U}(s;W_{j},W_{k})\,ds-\int_{0}^{1}S(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{H}(s;W_{j},W_{k})\,ds,$
then the numerical method can be written as (14)-(15)-(16) with
$D_{j,k}=A_{j,k}(U_{k}-U_{j})-S_{j,k}(H(x_{k})-H(x_{j})),$ (26)
with
$A_{j,k}=A_{\widetilde{\Psi}}(W_{j},W_{k}),\quad S_{j,k}=S_{\widetilde{\Psi}}(W_{j},W_{k}).$
Accordingly, if the LF splitting has been chosen, one has then
$D^{\pm}_{j,k}=A^{\pm}_{j,k}(U_{k}-U_{j})-S^{\pm}_{j,k}(H(x_{k})-H(x_{j})),$ (27)
where
$A^{\pm}_{j,k}=\frac{1}{2}\left(A_{j,k}\pm\alpha I\right),\quad S^{\pm}_{j,k}=\frac{1}{2}S_{j,k}.$
If the upwind approach is followed, the fluctuations are computed as follows: given two indices $j,k$, first the coordinates $\{\alpha_{j,k;l}\}_{l=1}^{N}$ of $U_{k}-U_{j}-A_{j,k}^{-1}S_{j,k}(H(x_{k})-H(x_{j}))$ in the basis of eigenvectors of the Roe matrix $A_{j,k}$ are computed:
$U_{k}-U_{j}-A_{j,k}^{-1}S_{j,k}(H(x_{k})-H(x_{j}))=\sum_{l=1}^{N}\alpha_{j,k;l}\vec{r}_{j,k;l}.$ (28)
Then, one has:
$D_{j,k}=A_{j,k}(U_{k}-U_{j})-S_{j,k}(H(x_{k})-H(x_{j}))=\sum_{i=1}^{N}\alpha_{j,k;l}\lambda_{j,k;l}\vec{r}_{j,k;l},$
and then
$D_{j,k}^{\pm}=\sum_{i=1}^{N}\alpha_{j,k;l}\lambda^{\pm}_{j,k;l}\vec{r}_{j,k;l}.$
The reconstruction is then performed as in the case of homogeneous problems.
The well-balanced property of methods desiged using this first strategy is given by the following result:
Proposition 2.
Let $U^{*}(x)$ be a stationary solution of system (24). Then, if the matrix splitting is such that
$D_{i,j}=0\implies D^{\pm}_{i,j}=0,\quad\text{for }j=i-k,\dots,i+k,$ (29)
and the chosen family of paths $\widetilde{\Psi}$ is such that, given $x_{L},x_{R}\in\mathbb{R}$ the functions
$s\in[0,1]\to\Psi_{U}\left(s;\left[\begin{array}[]{c}U^{*}(x_{L})\\ H(x_{L})\end{array}\right];\left[\begin{array}[]{c}U^{*}(x_{R})\\ H(x_{R})\end{array}\right]\right)$
and
$x\in[x_{L},x_{R}]\to U^{*}(x)$
define the same curve in $\Omega$, then the numerical method (14)-(15)-(16)-(25) is well-balanced for $U^{*}$, i.e. $\{U^{*}(x_{i})\}$ is an equilibrium of the ODE system (14).
Proof.
Let us check that $\{U^{*}(x_{i})\}$ is an equilibrium for the ODE system defined by the semi-discretized numerical method (14), i.e. that
$\widehat{D}_{i-1/2}^{+}+\widehat{D}_{i+1/2}^{-}=0,$
for $W_{i}=[U^{*}(x_{i}),H(x_{i})]^{T}$. In effect, we have
$D_{i,j}$ $=$ $\int_{0}^{1}A(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{U}(s;W_{j},W_{k})\,ds-\int_{0}^{1}S(\Psi_{U}(s;W_{j},W_{k}))\partial_{s}\Psi_{H}(s;W_{j},W_{k})\,ds$
[…]
$=$ $0,$
where a change of parameter has been applied to obtain the second equality and the fact that $U^{*}$ is a stationary solution has been used in the third one. Then, due to (29), from (15)-(16) we derive
$\widehat{D}_{i-1/2}^{+}=\widehat{D}_{i+1/2}^{-}=0,$
what ends the proof. ∎
Observe that (29) is satisfied for the upwind approach but it is not for the LF splitting is used: in effect, if $D_{i,j}=0$ one has
$D^{\pm}_{i,j}=\pm\frac{\alpha}{2}(U_{j}-U_{i})$
which is in general different from 0. Therefore, Strategy 1 for well-balancing cannot be combined with the choice of LF splitting.
3.2 Strategy 2
Strategy 2 extends to the nonconservative system the technique proposed in [38] for systems of balance laws: to obtain a fully well-balanced method, let us denote by $U^{*}_{i}$ the stationary solution satisfying
$U^{*}_{i}(x_{i})=U_{i},$
if it exists. Then, the equalities
$S(U_{i})H_{x}(x_{i})=S(U^{*}_{i}(x_{i}))H_{x}(x_{i})=A(U^{*}(x_{i}))(U^{*}_{i})_{x}(x_{i})$
can be used to approximate the source term, which leads to the approximate equation
$U_{t}+A(U)U_{x}-A(U^{*}_{i})(U^{*}_{i})_{x}(x_{i})=0,$
that is exactly at $x_{i}$. The idea is then to consider the numerical method (14) where now
$\widehat{D}^{-}_{i+1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i,i-k}-D^{*,+}_{i,i-k},\dots,D^{+}_{i,i+k}-D^{*,+}_{i,i+k})$ (30)
$+\mathcal{R}^{R}(D^{-}_{i,i+1-k}-D^{*,-}_{i,i+1-k},\dots,D^{-}_{i,i+1+k}-D^{*,-}_{i,i+1+k}),$
$\widehat{D}^{+}_{i-1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i-1-k,i}-D^{*,+}_{i-1-k,i},\dots,D^{+}_{i-1+k,i}-D^{*,+}_{i-1+k,i})$ (31)
$+\mathcal{R}^{R}(D^{-}_{i-k,i}-D^{*,-}_{i-k,i},\dots,D^{-}_{i+k,i}-D^{*,-}_{i+k,i}).$
with
$D^{*,\pm}_{j,k}=\int_{0}^{1}A^{\pm}(\Psi(s;U_{i}^{*}(x_{j}),U_{i}^{*}(x_{k}))\partial_{s}\Psi(s;U^{*}_{i}(x_{j}),U_{i}^{*}(x_{k}))\,ds,$ (32)
or
$D^{*,\pm}_{j,k}=A^{\pm}_{\Psi}(U_{i}^{*}(x_{j}),U_{i}^{*}(x_{k}))(U^{*}_{i}(x_{k})-U_{i}^{*}(x_{j})),$ (33)
if a Roe linearization of $A$ is available.
Proposition 3.
The numerical method (14)-(15)-(16) is fully well-balanced, i.e. $\{U^{*}(x_{i})\}$ is an equilibrium of the ODE system (14) for every stationary solution $U^{*}$.
Proof.
It is clear that
$U^{*}_{i}\equiv U^{*}$
is the stationary solution that satisfies
$U^{*}_{i}(x_{i})=U^{*}(x_{i})$
for all $i$. Therefore,
$D^{*,\pm}_{j,k}=D^{\pm}_{j,k}$
so that
$\widehat{D}^{+}_{i+1/2}=\widehat{D}^{+}_{i-1/2}=0,$
what ends the proof. ∎
Remark.
Observe that, to prove the result, it has been assumed that the matrix splitting used to approximate $A(U^{*}_{i})\partial_{x}U^{*}_{i}$ is exactly the same as the one used to approximate $A(U)U_{x}$. In particular, the same constant $\alpha$ has to be used in the LF splitting.
This strategy can be easily adapted to design numerical methods that only preserve a given $m$-parameter family of stationary solutions
$U^{*}(x;C_{1},\dots,C_{m})$
with $m<N$. The idea is to select the function of this family
$U^{*}_{i}(x)=U^{*}(x;C_{i,1},\dots,C_{i,m})$
whose parameters $C_{i,j}$, $j=1,\dots,m$ are such that
$u^{*}_{i;j_{l}}(x_{i})=u_{i;j_{l}},\quad l=1,\dots,m,$
where $u^{*}_{i;j}$ and $u_{i;j}$ represent respectively the $j$th component of $U^{*}_{i}(x)$ and $U_{i}$; and $j_{1},\dots,j_{m}$ are a set of pre-determined indices whose selection depends on the particular problems. The numerical method is in this case as follows:
$U_{i}^{\prime}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2}+\widehat{D}^{+}_{i-1/2}\right)=(S(U_{i})-S(U^{*}_{i}(x_{i})))H_{x}(x_{i}),$ (34)
together with (30)-(31)-(32). It can be trivially checked that it preserves the stationary solutions of the $m$-parameter family.
4 Extension to 2D systems
4.1 Homogeneous problems
This section extends the 1D path-conservative fifth-order WENO scheme to 2D nonconservative systems of the form
$U_{t}+A_{1}(U)U_{x}+A_{2}(U)U_{y}=0.$ (35)
The system is supposed again to be strictly hyperbolic, i.e. for all $U$ and all $\theta\in[0,2\pi)$, the matrix
$\cos(\theta)A_{1}(U)+\sin(\theta)A_{2}(U)$
has $N$ different real eigenvalues. Systems with flux terms and nonconservative products
$U_{t}+F(U)_{x}+G(U)_{y}+C(U)U_{x}+D(U)U_{y}=0,$ (36)
can be considered as particular cases in which $A_{1}(U)=J(F(U))+C(U),A_{2}(U)=J(G(U))+D(U).$
We consider uniform Cartesian meshes with points $(x_{i},y_{j})$ with steps $\Delta x$ and $\Delta y$ in the $x$ and $y$ direction. WENO methods can be extended dimension by dimension:
$U^{{}^{\prime}}_{i,j}=-\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2,j}+\widehat{D}^{+}_{i-1/2,j}\right)-\frac{1}{\Delta y}\left(\widehat{D}^{-}_{i,j+1/2}+\widehat{D}^{+}_{i,j-1/2}\right),$ (37)
where $U_{i,j}\approx U(x_{i},y_{j})$ and
$\widehat{D}^{-}_{i+1/2,j}$ $=$ $\mathcal{R}^{L}(D^{+}_{i,i-k;j},\dots,D^{+}_{i,i+k;j})+\mathcal{R}^{R}(D^{-}_{i,i+1-k;j},\dots,D^{-}_{i,i+1+k;j}),$ (38)
$\widehat{D}^{+}_{i-1/2,j}$ $=$ $\mathcal{R}^{L}(D^{+}_{i-1-k,i;j},\dots,D^{+}_{i-1+k,i;j})+\mathcal{R}^{R}(D^{-}_{i-k,i;j},\dots,D^{-}_{i+k,i;j}),$ (39)
$\widehat{D}^{-}_{i,j+1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i;j,j-k},\dots,D^{+}_{i;j,j+k})+\mathcal{R}^{R}(D^{-}_{i;j,j+1-k},\dots,D^{-}_{i;j,j+1+k}),$ (40)
$\widehat{D}^{+}_{i,j-1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i;j-1-k,j},\dots,D^{+}_{i;j-1+k,j})+\mathcal{R}^{R}(D^{-}_{i;j-k,j},\dots,D^{-}_{i;j+k,j}).$ (41)
Here, the following notation has been used:
$D^{\pm}_{i,k;j}$ $=$ $\int_{0}^{1}A_{1}^{\pm}(\Psi(s;U_{i,j},U_{k,j}))\partial_{s}\Psi(s;U_{i,j},U_{k,j})\,ds,$ (42)
$D^{\pm}_{i;j,k}$ $=$ $\int_{0}^{1}A_{2}^{\pm}(\Psi(s;U_{i,j},U_{i,k}))\partial_{s}\Psi(s;U_{i,j},U_{i,k})ds,$ (43)
where $A_{j}^{\pm}(U)$ represents the splitting chosen for $A_{j}(U)$: LF splitting can be readily extended to 2D and the upwind splitting as well if a Roe linearization is available. The extension of the strategy described in Section 2.1 to reconstruct characteristic variables is straightforward as well.
Proposition 1 can be easily extended to prove that (37) is a numerical method of order $p=2k+1$.
4.2 Problems with source terms
We now consider systems of the form
$U_{t}+A_{1}(U)U_{x}+A_{2}(U)U_{y}=S_{1}(U)H_{x}+S_{2}(U)H_{y},$ (44)
where $H(x,y)$ is again a known function. When a Roe linearization is available, Strategy 1 described in Section 3 can be readily extended to problem (44): it is enough to redefine the fluctuations as follows:
$D^{\pm}_{i,k;j}$ $=$ $A^{\pm}_{1;i,k;j}(U_{k,j}-U_{i,j})-S^{\pm}_{1;i,k;j}(H(x_{k},y_{j})-H(x_{i},y_{j})),$ (45)
$D^{\pm}_{i;j,k}$ $=$ $A^{\pm}_{2;i;j,k}(U_{i,k}-U_{i,j})-S^{\pm}_{2;i;j,k}(H(x_{i},y_{k})-H(x_{i},y_{j})),$ (46)
where $A_{1;i,k;j}$, $S_{1;i,k;j}$, represent respectively the intermediate matrix and source term given by the Roe linearization for the states $U_{i,j}$ and $U_{k,j}$ and $A_{2;i;j,k}$, $S_{2;i;j,k}$ the corresponding to the states $U_{i,j}$, $U_{i,k}$.
If (29) is assumed, it can be shown as in Proposition 2 that the numerical method is well-balanced for stationary solutions that satisfy that given $(x_{L},y),(x_{R},y),(x,y_{D}),(x,y_{U})\in\mathbb{R}^{2}$ the functions
$s\in[0,1]\to\Psi_{U}\left(s;\left[\begin{array}[]{c}U^{*}(x_{L},y)\\ H(x_{L},y)\end{array}\right];\left[\begin{array}[]{c}U^{*}(x_{R},y)\\ H(x_{R},y)\end{array}\right]\right),$ (47)
$s\in[0,1]\to\Psi_{U}\left(s;\left[\begin{array}[]{c}U^{*}(x,y_{D})\\ H(x,y_{U})\end{array}\right];\left[\begin{array}[]{c}U^{*}(x,y_{D})\\ H(x,y_{U})\end{array}\right]\right),$ (48)
define respectively the same curves in $\Omega$ as
$x\in[x_{L},x_{R}]\to U^{*}(x,y),$ (49)
$y\in[y_{D},y_{U}]\to U^{*}(x,y).$ (50)
This geometrical property is much more restrictive in 2D than in 1D and, in general, only stationary solutions that are essentially 1D or some particular families of stationary solutions satisfy them, as will be seen in the two-layer shallow-water case.
Strategy 2 described in Section 3.2 cannot be extended in general since, to implement it, a stationary solution $U^{*}_{i,j}$ satisfying
$U^{*}_{i,j}(x_{i},y_{j})=U_{i,j}$
is necessary, but now stationary solutions satisfy the PDE system
$A_{1}(U^{*})\partial_{x}U^{*}+A_{2}(U^{*})\partial_{y}U^{*}=S_{1}(U^{*})H_{x}+S_{2}(U^{*})H_{y},$
and one point value is not enough to determine one of them: in general, there will be infinitely many or none. Nevertheless, Strategy 2 can be extended easily to 2D problems to design numerical methods that preserve a given family of $m$-parameter stationary solutions
$U^{*}(x,y;C_{1},\dots,C_{m})$
with $m<N$. As in the 1D case, first the value of the parameters are imposed so that the value of $m$ pre-determined components at $(x_{i},y_{j})$ of the corresponding element of the family coincide with those of $U_{i,j}$. Once an element of the family $U^{*}_{i,j}$ has been determined, the numerical solution is written as follows:
$U^{{}^{\prime}}_{i,j}+\frac{1}{\Delta x}\left(\widehat{D}^{-}_{i+1/2,j}+\widehat{D}^{+}_{i-1/2,j}\right)+\frac{1}{\Delta y}\left(\widehat{D}^{-}_{i,j+1/2}+\widehat{D}^{+}_{i,j-1/2}\right)$ (51)
$\qquad\qquad=(S_{1}(U_{i,j})-S_{1}(U^{*}_{i,j}(x_{i},y_{j})))H_{x}(x_{i},y_{j})+(S_{2}(U_{i,j})-S_{2}(U^{*}_{i,j}(x_{i},y_{j})))H_{y}(x_{i},y_{j}),$
with
$\widehat{D}^{-}_{i+1/2,j}$ $=$ $\mathcal{R}^{L}(D^{+}_{i,i-k;j}-D^{*,+}_{i,i-k;j},\dots,D^{+}_{i,i+k;j}-D^{*,+}_{i,i+k;j})$ (52)
$+\mathcal{R}^{R}(D^{-}_{i,i+1-k;j}-D^{*,-}_{i,i+1-k;j},\dots,D^{-}_{i,i+1+k;j}-D^{*,-}_{i,i+1+k;j}),$
$\widehat{D}^{+}_{i-1/2,j}$ $=$ $\mathcal{R}^{L}(D^{+}_{i-1-k,i;j}-D^{*,+}_{i-1-k,i;j},\dots,D^{+}_{i-1+k,i;j}-D^{*,+}_{i-1+k,i;j})$ (53)
$+\mathcal{R}^{R}(D^{-}_{i-k,i;j}-D^{*,-}_{i-k,i;j},\dots,D^{-}_{i+k,i;j}-D^{*,-}_{i+k,i;j}),$
$\widehat{D}^{-}_{i,j+1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i;j,j-k}-D^{*,+}_{i;j,j-k},\dots,D^{+}_{i;j,j+k}-D^{*,+}_{i;j,j+k})$ (54)
$+\mathcal{R}^{R}(D^{-}_{i;j,j+1-k}-D^{*,-}_{i;j,j+1-k},\dots,D^{-}_{i;j,j+1+k}-D^{*,-}_{i;j,j+1+k}),$
$\widehat{D}^{+}_{i,j-1/2}$ $=$ $\mathcal{R}^{L}(D^{+}_{i;j-1-k,j}-D^{*,+}_{i;j-1-k,j},\dots,D^{+}_{i;j-1+k,j}-D^{*,+}_{i;j-1+k,j})$ (55)
$+\mathcal{R}^{R}(D^{-}_{i;j-k,j}-D^{*,-}_{i;j-k,j},\dots,D^{-}_{i;j+k,j}-D^{*,-}_{i;j+k,j}).$
Here, as in the 1D case, in the fluctuations $D^{*,\pm}_{k,l;m}$, $U_{k,m}$ and $U_{l,m}$ are replaced by $U^{*}_{i,j}(x_{k},y_{m})$ and $U^{*}_{i,j}(x_{l},y_{m})$, and in the fluctuations $D^{*,\pm}_{k;l,m}$, $U_{k,l}$ and $U_{k,m}$ are replaced by $U^{*}_{i,j}(x_{k},y_{l})$ and $U^{*}_{i,j}(x_{k},y_{m})$.
5 Application to the two-layer shallow water model
In this Section, we consider the 2D two-layer shallow water system that governs the flow of two superposed layers of immiscible fluids with different constant densities:
$\left(h_{1}\right)_{t}+\left(h_{1}u_{1,1}\right)_{x}+\left(h_{1}u_{1,2}\right)_{y}=0,$ (56)
$\left(h_{1}u_{1,1}\right)_{t}+\left(h_{1}u_{1,1}^{2}+\frac{g}{2}h_{1}^{2}\right)_{x}+\left(h_{1}u_{1,1}u_{1,2}\right)_{y}=-gh_{1}Z_{x}-gh_{1}\left(h_{2}\right)_{x},$
$\left(h_{1}u_{1,2}\right)_{t}+\left(h_{1}u_{1,1}u_{1,2}\right)_{x}+\left(h_{1}u_{1,2}^{2}+\frac{g}{2}h_{1}^{2}\right)_{y}=-gh_{1}Z_{y}-gh_{1}\left(h_{2}\right)_{y},$
$\left(h_{2}\right)_{t}+\left(h_{2}u_{2,1}\right)_{x}+\left(h_{2}u_{2,2}\right)_{y}=0,$
$\left(h_{2}u_{2,1}\right)_{t}+\left(h_{2}u_{2,1}^{2}+\frac{g}{2}h_{2}^{2}\right)_{x}+\left(h_{2}u_{2,1}u_{2,2}\right)_{y}=-gh_{2}Z_{x}-grh_{2}\left(h_{1}\right)_{x},$
$\left(h_{2}u_{2,2}\right)_{t}+\left(h_{2}u_{2,1}u_{2,2}\right)_{x}+\left(h_{2}u_{2,2}^{2}+\frac{g}{2}h_{2}^{2}\right)_{y}=-gh_{2}Z_{y}-grh_{2}\left(h_{1}\right)_{y},$
where
  • $h_{k}$, $k=1,2$ is the thickness of the $k$th layer;
  • $\vec{u}_{k}=(u_{k,1},u_{k,2})$, $k=1,2$ is the velocity of the $k$th layer;
  • $r=\frac{\rho_{1}}{\rho_{2}}$, where $\rho_{k}$, $k=1,2$ is the density of the $k$th layer ($\rho_{1}<\rho_{2}$). (index 1 corresponds to the upper layer);
  • $c_{k}=\sqrt{gh_{k}}$, $k=1,2$;
  • $Z(x,y)$ is the bottom function.
The equations can be written in the form (44) with $N=6$, $H\equiv-Z$,
$U=\left[\begin{array}[]{c}h_{1}\\ h_{1}u_{1,1}\\ h_{1}u_{1,2}\\ h_{2}\\ h_{2}u_{2,1}\\ h_{2}u_{2,2}\end{array}\right],$
$A_{1}(U)=\left[\begin{array}[]{cccccc}0&1&0&0&0&0\\ c_{1}^{2}-u_{1,1}^{2}&2u_{1,1}&0&c_{1}^{2}&0&0\\ -u_{1,1}u_{1,2}&u_{1,2}&u_{1,1}&0&0&0\\ 0&0&0&0&1&0\\ rc_{2}^{2}&0&0&c_{2}^{2}-u_{2,1}^{2}&2u_{2,1}&0\\ 0&0&0&-u_{2,1}u_{2,2}&u_{2,2}&u_{2,1}\end{array}\right],$ (57)
$A_{2}(U)=\left[\begin{array}[]{cccccc}0&0&1&0&0&0\\ -u_{1,1}u_{1,2}&u_{1,2}&u_{1,1}&0&0&0\\ c_{1}^{2}-u_{1,2}^{2}&0&2u_{1,2}&c_{1}^{2}&0&0\\ 0&0&0&0&0&1\\ 0&0&0&-u_{2,1}u_{2,2}&u_{2,2}&u_{2,1}\\ rc_{2}^{2}&0&0&c_{2}^{2}-u_{2,2}^{2}&0&2u_{2,2}\end{array}\right],$ (58)
$S_{1}(U)=\left[\begin{array}[]{c}0\\ gh_{1}\\ gh_{1}\\ 0\\ 0\\ 0\end{array}\right],\quad S_{2}(U)=\left[\begin{array}[]{c}0\\ 0\\ 0\\ 0\\ gh_{2}\\ gh_{2}\end{array}\right].$ (59)
The characteristic equation of $A_{1}(U)$ is
$(\lambda-u_{1,1})(\lambda-u_{2,1})\Bigl{(}\bigl{(}(\lambda-u_{1,1})^{2}-gh_{1}\bigr{)}\bigl{(}(\lambda-u_{2,1})^{2}-gh_{2}\bigr{)}-rg^{2}h_{1}h_{2}\Bigr{)}=0.$
The eigenvalues are then the four roots $\lambda_{k}$, $k=1,..,4$ of the equation
$\bigl{(}(\lambda-u_{1,1})^{2}-gh_{1}\bigr{)}\bigl{(}(\lambda-u_{2,1})^{2}-gh_{2}\bigr{)}=rg^{2}h_{1}h_{2},$
and
$\lambda_{5}=u_{1,1},\quad\lambda_{6}=u_{2,1}.$
The corresponding eigenvectors are
$\vec{R}_{k}=\left[\begin{array}[]{c}1\\ \lambda_{k}\\ u_{1,2}\\ \mu_{k}\\ \mu_{k}\lambda_{k}\\ \mu_{k}u_{2,2}\end{array}\right],\quad k=1,\dots,4,\quad\vec{R}_{5}=\left[\begin{array}[]{c}0\\ 0\\ 1\\ 0\\ 0\\ 0\end{array}\right],\quad\vec{R}_{6}=\left[\begin{array}[]{c}0\\ 0\\ 0\\ 0\\ 0\\ 1\end{array}\right],$ (60)
with
$\mu_{k}=\frac{(\lambda_{k}-u_{1,1})^{2}}{c_{1}^{2}}-1.$
The eigenvalues and eigenvectors of $A_{2}(U)$ can be computed similarly.
For $r\approx 1$, first-order approximation of $\lambda_{1},\dots,\lambda_{4}$ is given similar as [42]:
$\lambda_{ext}^{\pm}=U_{m}\pm\sqrt{g\left(h_{1}+h_{2}\right)},\quad\lambda_{\text{int }}^{\pm}=U_{c}\pm\sqrt{g^{\prime}\frac{h_{1}h_{2}}{h_{1}+h_{2}}\left[1-\frac{\left(u_{1}-u_{2}\right)^{2}}{g^{\prime}\left(h_{1}+h_{2}\right)}\right]},$ (61)
where
$U_{m}=\frac{h_{1}u_{1}+h_{2}u_{2}}{h_{1}+h_{2}},\quad U_{c}=\frac{h_{1}u_{2}+h_{2}u_{1}}{h_{1}+h_{2}},$
and $g^{\prime}=\left(1-r\right)g$ is the reduced gravity. Observe that $\lambda_{\text{int }}$ become complex when
$\frac{\left(u_{1}-u_{2}\right)^{2}}{g^{\prime}\left(h_{1}+h_{2}\right)}>1,$
so that the system is expected to lose hyperbolicity in these cases. Numerical techniques to overcome sporadic episodes of loss of hyperbolicity can be found in [10, 30].
The family of straight segments will be considered here to compute non-conservative products:
$\Psi(s;U_{L},U_{R})=U_{L}+s(U_{R}-U_{L}),\quad s\in[0,1].$
Steady-states solutions that correspond to water-at-rest equilibria constitute a 2-parameter family:
$u_{1,1}=u_{1,2}=u_{2,1}=u_{2,2}\equiv 0,\;h_{1}=C_{1},\;h_{2}=-Z+C_{2},$
where $C_{1}$, $C_{2}$ are arbitrary parameters ($C_{1}>0$, $C_{2}>\max(Z)$). We will focus here on methods that preserve this family.
The 1D two-layer shallow-water model will be also considered: it can be written in the form (24) with $N=4$, $H\equiv-Z$,
$U=\left[\begin{array}[]{c}h_{1}\\ h_{1}u_{1}\\ h_{2}\\ h_{2}u_{2}\end{array}\right],\quad A(U)=\left[\begin{array}[]{cccc}0&1&0&0\\ gh_{1}-u_{1}^{2}&2u_{1}&gh_{1}&0\\ 0&0&0&1\\ rgh_{2}&0&gh_{2}-u_{2}^{2}&2u_{2}\end{array}\right],\quad S=\left[\begin{array}[]{c}0\\ gh_{1}\\ 0\\ gh_{2}\end{array}\right],$ (62)
where now $u_{k}$, $h_{k}$, $k=1,2$ are respectively the velocity and thickness of the layers.
The following Roe linearization corresponding to the family of straight segments is available (see [25]):
$A(U_{L},U_{R})=\left[\begin{array}[]{ccccc}0&1&0&0\\ -\bar{u}_{1}^{2}+\bar{c}_{1}^{2}&2\bar{u}_{1}&\bar{c}_{1}^{2}&0\\ 0&0&0&1\\ r\bar{c}_{2}^{2}&0&-\bar{u}_{2}^{2}+\bar{c}_{2}^{2}&2\bar{u}_{2}\end{array}\right],\quad S(U,V)=\left[\begin{array}[]{c}0\\ g\bar{h}_{1}\\ 0\\ g\bar{h}_{2}\end{array}\right],$
where
$\bar{u}_{k}=\frac{\sqrt{h_{L,k}}u_{L,k}+\sqrt{h_{R,k}}u_{R,k}}{\sqrt{h_{L,k}}+\sqrt{h_{R,k}}},\quad\bar{h}_{k}=\frac{h_{L,k}+h_{R,k}}{2},\quad\bar{c}_{k}=\sqrt{g\bar{h}_{k}},\quad k=1,2.$
This Roe matrix and its natural extension to 2D will be used to implement WENO methods.
Both Strategies 1 and 2 described in Section 3 for 1D problems and 4.2 for 2D problems can be followed to design numerical methods that preserve water-at-rest solutions. In effect, for Strategy 1, the equality of the curves given by (47) and (49) or those given by (48) and (50) can be easily checked if the family of straight segments is chosen: it derives from the linear nature of the relationships between variables that characterize water-at-rest solutions. On the other hand, if Strategy 2 is followed, $U^{*}_{i,j}$ is chosen as follows;
$U^{*}_{i,j}(x,y)=\left[h_{1;i,j},0,0,-Z(x,y)+Z(x_{i},y_{j})+h_{2;i,j},0,0\right]^{T}.$
When the upwind-based splitting is used, systems of the form
$R\cdot\vec{\alpha}=U_{R}-U_{L}-A_{k}^{-1}S_{k}(H_{R}-H_{L}),$ (63)
have to be solved to compute $\alpha_{i}$, $i=1,\dots,6$ such that (28) is satisfied. Here, $A_{k}=A_{k}(U_{L},U_{R})$, $S_{k}=S_{k}(U_{L},U_{R})$, $k=1,2$ represent a Roe linearization in the $x$ or $y$ direction, and
$R=\left[\begin{array}[]{c|c|c}\vec{R}_{1}&\dots&\vec{R}_{6}\end{array}\right],\quad\vec{\alpha}=\left[\begin{array}[]{c}\alpha_{1}\\ \dots\\ \alpha_{6}\end{array}\right].$
Let us suppose that $k=1$. In this case, $\vec{R}_{i}$, $i=1,\dots,6$ are given by (60). It can be checked that the system can be solved as follows:
  1. 1.
    Solve system
    $R\left[\begin{array}[]{c}\alpha_{1}\\ \alpha_{2}\\ \alpha_{3}\\ \alpha_{4}\end{array}\right]=V_{R}-V_{L}-A^{-1}S(H_{R}-H_{L}),$
    where
    $V_{L}=\left[\begin{array}[]{c}h_{1,L}\\ h_{1,L}u_{1,1,L}\\ h_{2,L}\\ h_{2,L}u_{2,1,L}\end{array}\right],\quad V_{R}=\left[\begin{array}[]{c}h_{1,R}\\ h_{1,R}u_{1,1,R}\\ h_{2,R}\\ h_{2,R}u_{2,1,R}\end{array}\right],$
    $R=\left[\begin{array}[]{ccc}1&\dots&1\\ \lambda_{1}&\dots&\lambda_{4}\\ \mu_{1}&\dots&\mu_{4}\\ \mu_{1}\lambda_{1}&\dots&\mu_{4}\lambda_{4}\end{array}\right],\quad A=\left[\begin{array}[]{cccccc}0&1&0&0\\ \bar{c}_{1}^{2}-\bar{u}_{1,1}^{2}&2\bar{u}_{1,1}&\bar{c}_{1}^{2}&0\\ 0&0&0&1\\ r\bar{c}_{2}^{2}&0&\bar{c}_{2}^{2}-\bar{u}_{2,1}^{2}&2\bar{u}_{2,1}\end{array}\right],\quad S=\left[\begin{array}[]{c}0\\ g\bar{h}_{1}\\ 0\\ g\bar{h}_{2}\end{array}\right].$
  2. 2.
    If $\lambda_{i}\not=0$, $i=5,6$ compute
    $\alpha_{5}$ $=$ $\frac{F_{3}-u_{1,2}\sum_{j=1}^{4}\alpha_{j}}{\lambda_{5}},$
    $\alpha_{6}$ $=$ $\frac{F_{6}-u_{2,2}\sum_{j=1}^{4}\mu_{j}\alpha_{j}}{\lambda_{6}}.$
    Otherwise (which is the case in water-at-rest stationary solutions) define
    $\alpha_{5}=\alpha_{6}=0.$
In other words, the system to be solved in the 2D case reduce to those arising in 1D problems. Observe that, following this algorithm, the cases in which eigenvalues $\lambda_{i}$, $i=5,6$, vanish and thus $A_{k}$ cannot be inverted, the coefficients $\alpha_{j}$ can be still computed.
6 Numerical solutions
In this section, we present some numerical results for the 1D and 2D two-layer shallow water equations.
Two different WENO methods have been implemented and tested:
  • WENO methods with upwind splitting and Strategy 1 in Section 3.1 for the treatment of the source terms: these family of schemes will be called Methods 1.
  • WENO methods with LF splitting and Strategy 2 in Section 3.2 for the treatment of the source terms : these family of schemes will be called Methods 2.
For all the methods:
  • fifth-order WENO is used in all of the numerical tests but in accuracy test;
  • the third order TVD Runge-Kutta method in [45] is used for time stepping, which is a convex combination of forward Euler steps;
  • the approach in [30] is followed to deal with the episodes of loss of hyperbolicity;
  • the computation of eigenvalues and eigenvectors are computed in analytic form according to (61) and (60), respectively, and the LAPACK library is used to solve linear systems.
  • $CFL=0.45$ is used in all cases.
  • the free boundary conditions are imposed excepted in the steady-state solutions and accuracy test.
The numerical experiments have been implemented using FORTRAN $90$ compiled with the INTEL ifort compiler run in a Dell Precision workstation with 24 CPU cores and 128 gigabytes of memory.
In many numerical tests the results are comparable and, when this is the case, we only show the results obtained with Methods 2: we will only show results obtained with schemes of both families if there are significant differences or to compare them.
6.1 1D two-layer shallow water model
6.1.1 Small perturbations of a steady-state solution
We use this test, taken from [32], to verify the well-balanced property. The density ratio is $r=0.98$ and the gravitational constant is $g=10$. We consider smooth and discontinuous bottom topography. The smooth bottom topography is given by
$Z(x)=\begin{cases}0.25(\cos(10\pi(x-0.5))+1)-2,&\text{ if }0.4<x<0.6,\\ -2,&\text{ otherwise, }\end{cases}$
and the discontinuous one by
$Z(x)=\begin{cases}-1.5,&\text{ if }x>0.5,\\ -2,&\text{ otherwise. }\end{cases}$
The initial data is given by:
$\left(h_{1},h_{1}u_{1},h_{2},h_{2}u_{2}\right)=\begin{cases}(1+\sigma,0,-1-Z(x),0),&\text{ if }0.1<x<0.2,\\ (1,0,-1-Z(x),0),&\text{ otherwise. }\end{cases}$
The computational domain is $[-0.2,1]$ with extrapolation boundary conditions. We have first checked that for $\sigma=0$ the initial condition is preserved to machine accuracy, i.e. that the C-property is satisfied. Next, we set $\sigma=0.00001$. The numerical solutions computed with $200$ points are compared with a reference solution computed with $2000$ points. The final time is $T=0.15$. The numerical water surface, i.e. $\eta_{1}=-Z+h_{1}+h_{2}$, corresponding to the smooth and the discontinuous topographies are reported in Figure 1 (left and right respectively): it can be seen that small waves are captured without spurious oscillations as expected.
Figure 1: Section 6.1.1: small perturbation of a water-at-rest stationary solution with smooth (left) and discontinuous (right) bottom topographies. The numerical solution was obtained with 200 points and the reference solution was computed with 2000 points: water surface at $T=0.15$.
6.1.2 Riemann problems
We consider here two Riemann problems with flat bottom topography. This test is taken from [15]. In the first test, the computational domain is $[0,1]$ and the final time is $T=0.1$. The density ratio is $r=0.98$ and the gravitational constant is $g=10$. The initial condition is given by
$\left(h_{1},h_{1}u_{1},h_{2},h_{2}u_{2}\right)=\begin{cases}(0.50,1.250,0.50,1.250),&\text{ if }x<0.3,\\ (0.45,1.125,0.55,1.375),&\text{ otherwise. }\end{cases}$
The flat bottom is placed at $Z(x)=-1$. The water surface $E=-1+h_{1}+h_{2}$, a zoom of $h_{1}$, and the velocity $u_{1}$ obtained at the final time with a mesh of 200 cells are compared in Figure 2 with a reference solution obtained with 10000 points.
In the second test case, taken from [13], the computational domain is $[-5,5]$ and the final time is $T=1$. The density ratio is again $r=0.98$ and the gravitational constant is $g=9.81$. The initial data is given by
$\left(h_{1}(x,0),q_{1}(x,0),h_{2}(x,0),q_{2}(x,0)\right)=\begin{cases}(1.8,0,0.2,0),&\text{ if }x<0,\\ (0.2,0,1.8,0),&\text{ if }x>0.\end{cases}$
The bottom is placed now at $Z(x)=-2$. The water surface $E=-2+h_{1}+h_{2}$ and the interface $\eta=-2+h_{2}$ obtained with a 500-point mesh are shown in Figure 3, together with a reference solution obtained with a fine mesh with $5000$ points.
Figure 2: Section 6.1.2: Riemann problem 1. The numerical solution computed with 200 points is compared to a reference solution computed with 10000 points. Top left: water surface; top right: $h_{1}$ (zoom); bottom: $u_{1}$.
Figure 3: Section 6.1.2: Riemann problem 2. The numerical solution obtained with $500$ points is compared to a reference solution computed with $5000$ points. Left: water surface; right: interface.
6.1.3 An internal dam-break with flat bottom
In this test, taken from [25], an internal dam-break over a flat bottom is simulated. The computational domain is $[0,10]$ with free boundary conditions. The final time is $T=10$. The gravitational constant $g=10$ and the density ratio $r=0.98$. The initial condition is given by:
$\left(h_{1}(x,0),q_{1}(x,0),h_{2}(x,0),q_{2}(x,0)\right)=\begin{cases}(0.2,0,0.8,0),&\text{ if }x<5,\\ (0.8,0,0.2,0),&\text{ otherwise}.\end{cases}$
The results computed with a $200$-point mesh are compared in Figure 4 with a reference solution computed using $3200$ points: that solutions agree well with those in [25, 30].
Figure 4: Section 6.1.3: internal dam-break with flat bottom. The numerical solution computed with $200$ points is compared to a reference solution computed using $3200$ points: free surface (top-left), $h_{2}$ (top-right), $h_{1}u_{1}$ (down-left), and $h_{2}u_{2}$ (down-right) at time $T=10$.
6.1.4 Internal dam-break with non-flat bottom
In this test, taken from [22], a discontinuous stationary solution is reached starting from an internal dam-break initial condition given by
$\left(h_{1},q_{1},h_{2}+Z(x),q_{2}\right)(x,0)=\begin{cases}(1.6,0,-1.6,0)&\text{ if }x<0,\\ (0.7,0,-0.7,0)&\text{ otherwise, }\end{cases}$
and the bottom topography is given by
$Z(x)=0.25e^{-x^{2}}-2.$
The constant gravitational acceleration is $g=9.81$ and the density ratio is $r=0.998$. The computational domain is $[-5,5]$. The steady state obtained with $500$ points is shown in Figure 5. The converged results agree well with areference solutions computed with $2000$ grids and with those in [22].
Figure 5: Section 6.1.4: internal dam-break with non-flat bottom. The numerical solution obtained with $500$ points is compared to a reference solution computed with $2000$ points. Free surface, interface, and bottom topography: initial condition (left), reached steady state (right).
6.1.5 Internal dam break with loss of hyperbolicity
This test is taken from [24]. The constant density ratio is $r=0.98$ and the gravitational constant is $g=9.8$. The computational domain is $[0,10]$. The final time is $T=20$. The bottom topography and the initial condition are given now by
$Z(x)=0.47e^{-(x-5)^{2}},$
$h_{2}=\left\{\begin{array}[]{ll}0.5-Z(x),&\text{ if }x<5,\\ 0.97-Z(x),&\text{ otherwise, }\end{array}\quad h_{1}=1-h_{2}-Z(x),\quad u_{1}=u_{2}=0.\right.$
In this test, Methods 1 and 2 give different solutions due to the corrections of hyperbolicity losses: Figure 6 shows the numerical results computed with Methods 1 and 2 using $200$ points. The points in which a hyperbolicity loss has been detected are marked in purple.
Figure 6: Section 6.1.5: internal dam break with loss of hyperbolicity at $T=20$. The contrast between Method 1 and Method 2. Left: Method 1, Right: Method 2.
It can be observed that fewer corrections are needed with Method 1 (4 points) than with Method 2 (7 points). This is so during the whole simulation, what seems to indicate that Method 1 has better properties of hyperbolicity preservation.
6.1.6 Accuracy test
In this test, taken from [22], we check the empirical order of accuracy of the methods using a smooth solution. The bottom is flat, $g=9.81,r=0.98$, and the initial condition is given by:
$\left(h_{1}(x,0),q_{1}(x,0),h_{2}(x,0),q_{2}(x,0)\right)=\left(1-\frac{1}{2}\sin(8x),0,0.6+\frac{1}{2}\sin(8x),0\right).$
The computational domain is $[-\pi/2,3\pi/2]$ and periodic boundary conditions are considered. We compute the numerical solutions until $T=0.1$. We use 3rd, 5th, and 7th order WENO with linear weights for this smooth problem to check the numerical accuracy. The reference solutions for 3rd and 5th are computed by each order WENO scheme with 6400 uniform grids respectively. We use 12800 grids to obtain the reference solution for the 7th order accuracy test to get a more accurate reference. The third-order TVD Runge Kutta is used in all cases for time stepping and, in order to reach the same order of accuracy in time and space, the time-step is set to $\Delta t^{\frac{k}{3}}$, where $k$ is the WENO order and $\Delta t$ is the step given by the CFL condition.
Tables 1 and 2 show the $L^{\infty}$ error and the empirical order of convergence: in all cases, the optimal order is reached.
Table 1: Section 6.1.6. $L^{\infty}$ error and empirical order of convergence using WENO3 & WENO7.
N 3rd order test 7th order test
$h_{1}$ error order $h_{1}+h_{2}$ error order $h_{1}$ error order $h_{1}+h_{2}$ error order
25 3.32e-1 5.84e-3 1.62e-1 3.37e-3
50 1.10e-2 1.60 2.10e-3 1.48 1.20e-2 3.76 6.26e-4 2.43
100 2.08e-2 2.40 4.96e-4 2.08 9.58e-4 3.65 1.82e-5 5.10
200 3.19e-3 2.71 7.43e-5 2.74 1.94e-5 5.63 1.82e-7 6.64
400 4.24e-4 2.91 9.75e-6 2.93 2.04e-7 6.57 1.49e-9 6.94
800 5.31e-5 3.00 1.23e-6 2.99 1.64e-9 6.96 1.01e-11 7.20
Table 2: Section 6.1.6. Errors in $L^{\infty}$ norm and empirical order of convergence for WENO5.
N $h_{1}$ error order $h_{1}+h_{2}$ error order
25 5.38e-2 1.42e-3
50 4.65e-3 3.53 4.50e-4 1.66
100 4.84e-4 3.26 1.42e-5 4.98
200 1.95e-5 4.63 5.52e-7 4.69
400 5.92e-7 5.04 1.74e-8 4.99
800 1.76e-8 5.07 4.33e-10 5.33
6.2 2D two-layer shallow water model
6.2.1 2-D steady-state solution
This test is used to check the well-balanced property of the methods for water-at-rest solutions. The constant density ratio is $r=0.98$ and the gravitational constant is $g=10$. The bottom topography is given by a smooth function
$Z(x,y)=0.05e^{-100\left(x^{2}+y^{2}\right)}-1,$
and the initial condition is given by
$h_{1}=0.5,\;h_{2}=1-Z(x,y),\;u_{1,1}=u_{1,2}=u_{2,1}=u_{2,2}=0.$
The computational domain is $[-1,1]\times[-1,1]$. The final time is $T=0.1$. This initial boundary value problem is a modification of the exact C-property test, proposed in [46] for the shallow water equations. The numerical $L^{1}$ errors corresponding to two meshes of $50\times 50$ and $100\times 100$ points are shown in Table 3, as it can be seen the initial condition is preserved to machine accuracy.
We have used this test to compare the computational costs of Methods 1 and 2: the CPU times corresponding to a computation with $200\times 200$ points are shown in Table 4. As expected, Method 1 is more expensive, the speedup ratio being of around 1.3. Similar speedups are observed in other test cases.
Table 3: Section 6.2.1. $L^{1}$ errors at time $0.1$ using two meshes of $50\times 50$ and $100\times 100$ points.
$h_{1}$ $h_{1}u_{1}$ $h_{1}v_{1}$ $h_{2}$ $h_{2}u_{1}$ $h_{2}v_{2}$
$50\times 50$ 7.77e-16 1.36e-16 2.79e-16 9.81e-16 9.78e-15 1.36e-16
$100\times 100$ 1.98e-15 1.88e-16 5.71e-16 1.44e-15 1.32e-14 1.88e-16
Table 4: Section 6.2.1. Methods 1 and 2: CPU times and speedup.
CPU time Speedup ratio
Method 1 84.97 s
Method 2 65.44 s 1.3
6.2.2 Interface propagation with flat bottom
This test, taken from [31], is aimed to verify the robustness of the numerical method. The initial condition is
$\left(h_{1},h_{1}u_{1},h_{1}v_{1},h_{2},h_{2}u_{2},h_{2}v_{2}\right)(x,y,0)=\begin{cases}(0.50,1.250,1.250,0.50,1.250,1.250),&\text{ if }(x,y)\in\Omega,\\ (0.45,1.125,1.125,0.55,1.375,1.375),&\text{ otherwise,}\end{cases}$
where $\Omega=\{x<-0.5,y<0\}\cup\left\{(x+0.5)^{2}+(y+0.5)^{2}<0.25\right\}\cup\{x<0,y<-0.5\}$. The constant gravitational acceleration is $g=10$ and the density ratio is $r=0.98$. The flat bottom topography is given by $Z(x,y)\equiv-1$. The computational domain is $[-1,1]\times[-1,1]$ and the final time, $T=0.1$. The computational results obtained with two meshes of $400\times 400$ and $800\times 800$ are shown in Figure 7.
Figure 7: Section 6.2.2: interface propagation in 2D with flat bottom. Numerical solutions obtained with $400\times 400$ (top row) and $800\times 800$ (bottom row) grids: upper layer thickness $h_{1}$ (left) and water surface $h_{1}+h_{2}$ (right).
6.2.3 Interface propagation with non-flat bottom
This test is similar to the one in Section 6.2.2 but with a non-flat bottom given by
$Z(x,y)=0.05e^{-100\left(x^{2}+y^{2}\right)}-1,$
and the following initial data:
$\left(h_{1},h_{1}u_{1},h_{1}v_{1},h_{2},h_{2}u_{2},h_{2}v_{2}\right)(x,y,0)=\begin{cases}(0.50,1.250,1.250,0.50-Z,1.250,1.250),&\text{ if }(x,y)\in\Omega,\\ (0.45,1.125,1.125,0.55-Z,1.375,1.375),&\text{ otherwise. }\end{cases}$
The computational domain is $[-1,1]\times[-1,1]$ and the values of $g$ and $r$, and the final time are the same. We show again the results computed with $400\times 400$ and $800\times 800$ points in Figure 8. The numerical solutions are in good agreement with those presented in [33] but finer structures of the flow are actually captured here.
Figure 8: Section 6.2.3: interface propagation in 2-D with nonflat bottom. Numerical solutions computed with $400\times 400$ (top row) and $800\times 800$ (bottom row) grids: water surface $h_{1}+h_{2}+Z$ (left) and upper layer thickness $h_{1}$ (right).
6.2.4 Internal circular dam break over flat bottom topography
This example is taken from [9]: an internal circular dam-break problem with a flat bottom is considered. The initial conditions are given by
$\left(h_{1},q_{1},p_{1},h_{2},q_{2},p_{2}\right)(x,y,0)=\begin{cases}(1.8,0,0,-1.8-Z,0,0),&\text{ if }x^{2}+y^{2}>4,\\ (0.2,0,0,-0.2-Z,0,0),&\text{ otherwise. }\end{cases}$
In this test, we consider $g=9.81$, $r=0.998$, and $Z\equiv-2$. The final time is $T=20$.
Figure 9: Section 6.2.4: internal circular dam breaking in 2D with a flat bottom. Numerical solutions computed with $200\times 200$ grids. Diagonal slices of the water surface $h_{1}+h_{2}+Z$, interface $h_{2}+Z$ and bottom at times $t=4,6,10,14,16,20$ (from top to down and from left to right).
Diagonal slices $y=x$ of the numerical results computed with $200\times 200$ points for times $t=4,6,10,14,16,20$ are shown in Figure 9. We show the contour lines of the water interface for times $t=4,20$ in Figure 10. It is worth mentioning that the results agree well with those in [9, 22].
Figure 10: Section 6.2.4: internal circular dam breaking in 2D with a flat bottom. Numerical solutions computed with $200\times 200$ grids. Contour lines of the interface $h_{2}+Z$ at times $t=4$(left), $t=20$(right).
6.2.5 Internal circular dam break over nonflat bottom topography
In this test, we consider nonflat bottom
$Z(x,y)=0.5e^{\left(x^{2}+y^{2}\right)}-2,$
and the initial condition
$\left(h_{1},q_{1},p_{1},h_{2},q_{2},p_{2}\right)(x,y,0)=\begin{cases}(1.8,0,0,-1.8-Z,0,0),&\text{ if }x^{2}+y^{2}>1,\\ (0.2,0,0,-0.2-Z,0,0),&\text{ otherwise. }\end{cases}$
The constant gravitational acceleration is $g=9.81$ and the density ratio is $r=0.98$. The computational domain is $[-2,2]\times[-2,2]$ and a uniform mesh with $200\times 200$ grids is considered. The contour lines and diagonal slices of the water interface at $t=1,t=2$ are shown in Figure 11.
Figure 11: Section 6.2.5: internal circular dam break over nonflat bottom topography. Numerical solutions computed with $200\times 200$ grids. Contour lines of the interface $h_{2}+Z$ at times $t=1$(top left) and $t=2$(top right); diagonal slices of the interface at time $t=1$ (down left) and $t=2$ (down right).
7 Conclusion
A new family of high-order WENO finite-difference methods to solve hyperbolic nonconserative PDE systems have been proposed. These methods are based in a general strategy in which, instead of reconstructing fluxes using a WENO operator, what we reconstruct is the nonconservative products of the system which are computed using the selected family of paths. Moreover, if a Roe linearization is available, the nonconservative products can be computed through matrix-vector operations instead of path-integrals. The main advantages of this theory are the following:
  • the high-order accuracy in space only depends on the order of the selected WENO operator and no extra order-dependent correction terms are required;
  • no integrals in the cells have to be computed so that reconstructions with uniform accuracy in the entire cells are required.
The methods have been extended to systems with source terms and two strategies have been introduced to design well-balanced methods. This methodology has been successfully applied to obtain high-order numerical methods for the 1D and 2D two-layer shallow water equations that preserve water-at-rest steady states. A number of numerical tests confirm the high-order accuracy of the methdos as well as their shock-capturing and well-balanced properties.
The second-strategy introduced here to design well-balanced methods can be combined with the technique developed in [27] to obtain numerical methods that preserve both water-at-rest and moving equilibria for the 1D shallow-water model or some particular families of stationary solutions in the 2D case: this will be done in an upcoming work.
Another further development concerns the convergence of the methods: as it happens for general finite-difference and related methods, the convergence of the numerical results to functions that are weak solutions of the system according to the selected family of paths is not ensured for the schemes introduced here. Some techniques recently developed in the context of high-order finite-volume methods in [40, 41] can be adapted to the methods introduced here to ensure that isolated shocks that satisfy the generalized Rankine-Hugoniot associated to the selected family of paths are correctly captured.
Acknowledgements
The work of B. Ren is supported by the China Scholarship Council. The work of C. Parés is supported by Spanish projects RTI2018-096064-B-C1 and PID2022-137637NB-C21 funded by MCIN/AEI/10.13039/501100011033 and FSE+.